OpenStudy (karolynxx):

Can somebody help me learn how to solve these equations ? Solve the equation. y + 3 = –y + 9 a. y = 1 b. y = 3 c. y = 6 d. y = 9

OpenStudy (karolynxx):

help? @welshfella

OpenStudy (welshfella):

you need to get the y's to one side of the equation and the numbers on the other side so first add y to both sides y + y + 3 = -y + y + 9 now simplify 2y + 3 = 9 Next subtract 3 from both sides 2y + 3 - 3 = 9 - 3 2y = 6 finally divide both sides by 2 2y / 2 = 6/2 y = 3

OpenStudy (karolynxx):

Thank you!